(2n+1) C 0 + (2n+1) C 1 + (2n+1) C 2 +........+ (2n+1) C n = 2 x What is x?

(2n+1)C(2n+1)C(2n+1)C+........+ (2n+1)C= 2x
 
What is x? 

Grade:12

2 Answers

kanithi premsai
6 years ago
We know that, Nco+Nc1+......+Ncn=2 to the power N
Therefore the above sum of coefficients​ is 2to the power of 2n+1
 Therefore,
                    2x=2 to the power of 2n+1
                   : x= 2 to the power of 2n+1/2
                   : x=2 to the power of 2n
Therefore the value is 2 power 2n
Akshat Agarwal
6 years ago
Objectively u can put n=13C0 + 3C1= 1+3=4 x=2Similarly you can crosscheck the answer by putting n=2,3,4.........